FOURIER TRANSFORM UV EMISSION SPECTROSCOPY OF THE B ^{2}Σ^{+} − X ^{2}Σ^{+} BAND OF THE PN^{+}

\underline{TAKASHI IMAJO}, KEN TOKIEDA, KEIICHI TANAKA, and TAKEHIKO TANAKA, Department of Chemistry, Faculty of Science, Kyushu University 33, Hakozaki, Higashi ku, Fukuoka 812-8581, Japan..

High resolution Fourier transform emission spectrum of the B ^{2}Σ^{+} − X ^{2}Σ^{+} band of the PN^{+} ion was observed in the ultra violet region around 31000 cm^{−1} with Bruker IFS120HR. The rotationally resolved spectrum of PN^{+}, produced by an ac discharge of the PCl_{3}, N_{2}, and He gas mixture, was recorded for about 11 h accumulating the interferogram 1600 times. The monochromator placed just before a photomultiplier worked effectively as an optical filter with a band width of 200 cm^{−1}. The (0,0) band lines split into doublet due to the spin-rotation splitting were assigned to the R- and P- branch lines of N” up to 35. The determined band origin ν_{0}=31058.9963(81) cm^{−1} is consistent with the previous low resolution result [1]. The rotational, centrifugal distortion constants for both states were determined as well as the spin-rotation constant for the ground state. The B ^{2}Σ^{+} state has weaker bond than that of the ground state, since the bond length of the B ^{2}Σ^{+} (v=0) state (1.570 Å) is by 0.076 Å longer, and the force constant (k = 2.90 mdyne/Å) is by about 70 % smaller than those of the ground state. Present results support the ab initio calculation which predicts a shallow minimum for the B ^{2}Σ^{+} state due to avoided crossing [2].

1. I. K. Ahmad and P. A. Hamilton, J. Mol. Spectrosc., 163, 214 (1994).

2. F. Grein, Chem. Phys. Lett., 120, 383 (1988).